### v1

# Model-driven Stochastic Trace Clustering

Abstract
Process discovery algorithms automatically extract process models from event logs, but high variability often results in complex and hard-to-understand models. To mitigate this issue, trace clustering techniques group process executions into clusters, each represented by a simpler and more understandable process model. Model-driven trace clustering improves on this by assigning traces to clusters based on their conformity to cluster-specific process models. However, most existing clustering techniques rely on either no process model discovery, or non-stochastic models, neglecting the frequency or probability of activities and transitions, thereby limiting their capability to capture real-world execution dynamics. We propose a novel model-driven trace clustering method that optimizes stochastic process models within each cluster. Our approach uses entropic relevance, a stochastic conformance metric based on directly-follows probabilities, to guide trace assignment. This allows clustering decisions to consider both structural alignment with a cluster’s process model and the likelihood that a trace originates from a given stochastic process model. The method is computationally efficient, scales linearly with input size, and improves model interpretability by producing clusters with clearer control-flow patterns. Extensive experiments on public real-life datasets show that our method outperforms existing alternatives in representing process behavior and reveals how clustering performance rankings can shift when stochasticity is considered.
keywords:
Process Mining , Trace Clustering , Stochastic Models , Entropic Relevance
journal: Information Systems
[…]
[kul]organization=Research Center for Information Systems Engineering (LIRIS), KU Leuven, city=Leuven, country=Belgium
1 Introduction
Cluster analysis is a fundamental technique in data mining and machine learning, aimed at grouping entities such that those within the same cluster are more similar to each other than to those in different clusters. In process mining, trace clustering applies this principle to event logs by grouping process executions (cases) into smaller, more coherent subsets. This is particularly valuable for complex processes, as it enables the discovery of simpler, more interpretable process models for each cluster. Such models enhance the readability of discovered process models and can also improve the accuracy of conformance checking [1, 2, 3]. However, many existing trace clustering methods are unsupervised and optimize generic similarity objectives that are not explicitly aligned with improving process model quality or interpretability [1]. Model-driven trace clustering addresses this by iteratively discovering a process model for each cluster and assigning traces based on how well they conform to these models [1, 4].
In parallel, the emergence of the stochastic process mining paradigm has highlighted the importance of capturing the stochastic nature of real-world processes. Unlike purely structural models, stochastic process models account for the likelihood of behaviors, making them more suitable for applications such as simulation, prediction, and representative modeling of observed behavior.
Despite these developments, existing model-driven clustering methods have yet to incorporate the stochastic characteristics of event logs. To our knowledge, no approach currently integrates stochasticity into the clustering process to produce stochastic process models per cluster. To address this gap, we introduce a novel model-driven trace clustering method that incorporates stochastic conformance measuring into clustering decisions. Our method leverages entropic relevance (ER) [5], a stochastic conformance measure that quantifies the compression cost of traces in an event log based on directly follows probabilities. By minimizing ER within each cluster, the algorithm optimizes the representativeness of that cluster’s Directly-Follows Graph (DFG)—one of the most widely used model types in (commercial) process mining tools [6].
Our approach is computationally efficient, scaling linearly with the size of the event log, and supports integration with edge frequency filtering to mitigate the “spaghetti model” problem, potentially still present in downstream clusters. By promoting clusters with shared, high-probability directly-follows relations, it enhances model clarity and interpretability. Through extensive experiments on publicly available real-world datasets, we demonstrate that our method effectively captures control-flow behavior and yields interpretable, representative stochastic models. It also reveals that incorporating stochasticity can significantly alter the relative performance of clustering algorithms, underlining the importance of probabilistic considerations in trace clustering.
To this end, our work makes the following key contributions:
  • We introduce stochastic conformance-based model-driven trace clustering, explicitly incorporating probabilistic behavior in clustering decisions.
  • We propose an efficient algorithm using entropic relevance, optimizing within cluster process model coherence while maintaining scalability.
  • We provide a comprehensive evaluation framework, including diverse metrics to assess the impact of stochasticity on clustering rankings and process model interpretability.
Our work is directly relevant to industry applications, particularly DFG discovery tools, which frequently rely on cluster-like representations to extract common and exceptional behaviors. Given its low computational complexity, our method provides a practical extension to existing process mining visualization and conformance checking techniques for large-scale event logs.
The remainder of this paper is structured as follows: Section 2 introduces key terminology and discusses relevant background literature. Section 3 explains the workings of the entropic clustering algorithm. The experimental setup is detailed in Section 4, followed by the results presented in Section 5. Section 6 discusses these results along with potential limitations of the current approach. Finally, Section 7 provides a conclusion and outlines possible directions for future research. The code and the full results can be found online [^1] https://github.com/jaripeeperkorn/EntropicClustering.
2 Background
In this section, we provide an overview of process mining and its subdomain of trace clustering.
2.1 Event data and traces
The main data source used in process mining exists in the form of event logs, which record the executions of different activities within a business process. Each event typically includes a case ID that identifies the specific process instance it belongs to, an activity type that indicates what action was performed, and a timestamp or another attribute that defines the event’s position in the sequence. The ordered set of events for a case is referred to as a trace. Since the same sequence of activities may occur in multiple cases, an event log is formally a multiset of traces. In this work, we focus solely on control-flow information, meaning we represent traces as sequences of activity labels, disregarding other event attributes such as resources or timestamps. The set of all unique traces observed in an event log is referred to as the set of (control-flow) variants, with each variant’s frequency denoted by its multiplicity. A variant log is an abstraction of the full event log, capturing all distinct variants and their counts. To analyze behavior, traces are often compared to a process model such as a Petri net, which graphically depicts the logical structure and dynamics of a business process.
2.2 Process discovery & conformance checking
Process models can be manually designed or automatically discovered from event logs through process model discovery techniques, which aim to generate models that reflect and generalize the observed behavior. Well-known methods for automatically discovering Petri Nets include Heuristic Miner [7] and Inductive Miner [8], while other methods focus on other model types such as Fuzzy Miner [9] that can be used to create graph representations of the process at different levels of detail. Conformance checking techniques are used to compare an event log with a given process model, along different dimensions. Fitness measures how well the behavior found in the event log is described by the model, while precision measures how much of the allowed behavior is actually present in the event log. Typical fitness and precision metric include token-replay [10] and alignments [11, 12]. Generalization metrics aim to measure the model’s usefulness on unseen (but correct) process behavior. Simplicity reflects the readability, understandability, and complexity of the process, measured for example by the number of arcs [13]. Stochastic process models extend traditional process models by incorporating probabilities (or other quantitative aspects) of process executions. They provide insights into the likelihood of different behaviors. Stochastic conformance checking techniques incorporate these probabilities, e.g., by analyzing the statistical properties of execution patterns [14, 15, 16]. A widely used model type is the directly-follows graph (DFG), where nodes correspond to activity types and directed edges represent directly-follows (DF) relations observed in the log. DFGs are commonly used in commercial tools to visualize the process executions found in an event log [6]. By normalizing the edge frequencies with respect to the counts of their source activities, DFGs can also be interpreted as probabilistic (stochastic) models, with each DF relation assigned a transition probability. This property allows us to treat DFGs as stochastic process models in our work.
2.3 Trace clustering
Trace clustering refers to the task of grouping similar traces together based on control-flow patterns or other attributes. Clustering can help uncover subprocesses or behavioral variants, facilitating the discovery of simpler and more interpretable models for each group. A common line of work relies on vector-based methods, which transform traces into numerical feature vectors and apply standard clustering algorithms [2, 3, 17]. Context-aware approaches extend this by incorporating additional information about the execution context or structural dependencies [18]. A more process-centric alternative is provided by model-based clustering techniques. Ferreira et al. [4] proposed a method that learns a mixture of first-order Markov models using the Expectation-Maximization algorithm, but its computational cost limits scalability. ActiTraC [1] is another model-based approach that iteratively mines models using the Heuristic Miner algorithm [7] and evaluates the assignment of traces to clusters using fitness scores. Other works make use of alignments to perform trace clustering in different ways as well: using local alignments in combination with k-means clustering [19], by using model traces as centroids [20], or by explicitly using an available process model [21]. Despite existing advances, no method has yet integrated stochastic conformance information into the clustering process. In particular, model-driven clustering approaches have not leveraged the probabilistic behavior captured by stochastic process models. This gap motivates our contribution: a trace clustering technique that uses entropic relevance to evaluate and guide trace assignments.
3 Model-driven stochastic clustering
This section presents the foundations and implementation of our proposed approach, including the entropic relevance metric, the entropic clustering algorithm, and the initialization strategies used.
3.1 Entropic Relevance
We use entropic relevance (or ER), a stochastic conformance checking measure that quantifies the information-theoretic compression cost of traces in an event log, given the structure and likelihoods of a stochastic process model [5]. A lower compression cost (i.e., a lower ER score) indicates a better fit between the model and the event log. For each trace, the ER contribution is computed using the probability $p$ assigned to that trace by the model, calculated as: $-\text{log}_{2}(p)$. The total ER score for a given log-model combination is obtained by summing these contributions over all traces in the log. For non-fitting traces, a so-called background model is used, giving a compression cost that is completely based on the log only (and is generally much higher). ER gives a score that measures quality along both the fitness and precision dimension. By punishing for a lack of fitness through the background model, or from a stochastic point of view, lower values for $p$ are assigned to highly unlikely cases. If the model allows too much behavior (with high probability), i.e., it is imprecise, the probabilities of the fitting cases will drop as well, leading to an increase in ER score.
In our approach, we use DFGs, which can be converted to stochastic process models. The discovery of a DFG can be done fast and efficiently, by just counting how often each activity appears and the frequency with which one activity directly follows another in the traces of the event log. For our purposes, the DFG is stored as two dictionaries: one containing the activity types (nodes) and their counts, and one the directly-follows relations (edges) and their counts. Given a DFG, the probability of a specific trace is computed as the product of the transition (edge) probabilities along the trace. Each transition probability is calculated by dividing the count of the corresponding edge by the count of its source node. In the original ER implementation, stochastic process models are converted into stochastic deterministic finite automata (SDFAs) to compute trace probabilities. To replicate this in the DFG-based approach, special Beginning-Of-Sequence (BOS) and End-Of-Sequence (EOS) activities are introduced to all traces and included in the DFG. Incorporating these artificial transitions effectively replicates the initialization and termination probabilities that would otherwise be handled in an SDFA. The traces with the added BOS and EOS activities are also the ones used to discover the DFG, i.e., the DFGs will also include those nodes. Since all traces now start with BOS and end with EOS, the transition probabilities from the BOS node to any other node can be seen as initialization probabilities, and those from each node to EOS as termination probabilities. Our (simplified) ER calculation is formalized in Algorithm 1. In order to be able to compare ER scores across different clusters (or event logs) with different amounts of traces, we also opted to use the average ER score over traces, rather than the sum over all traces.
Algorithm 1 Compute (simplified) Entropic Relevance (ER) for a DFG-based Model
0: Variant log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$: $K$ distinct variants $t_{i}$, each with multiplicity $m_{i}$ directly-follows graph $G=(N,E,c_{N},c_{E})$: with $N=$ set of activity labels incl. BOS and $E=$ set of directed edges $c_{N}$ and $c_{E}$ the count of occurrences of the nodes and edges (the DFG). underflow threshold $\varepsilon$
0: Average $\mathrm{ER}$ score
1:$T\leftarrow\sum_{i=1}^{K}m_{i}$Total number of trace occurrences}
2:$\mathrm{ER}_{\text{sum}}\leftarrow 0$Sum of per-trace information costs}
3:for all $(\alpha,\beta)\in E$ do
4:    $P(\beta\mid\alpha)\leftarrow c_{E}(\alpha,\beta)/c_{N}(\alpha)$Probability of transition $\alpha\to\beta$}
5:end for
6:for $i=1$ to $K$ do
7:    Let $t_{i}=\langle a_{0},a_{1},\dots,a_{\ell}\rangle$$a_{0}=\mathrm{BOS},\,a_{\ell}=\mathrm{EOS}$}
8:    $p\leftarrow 1$Initialize probability of $t_{i}$}
9:    for $j=1$ to $\ell$ do
10:       $p\leftarrow p\times P(a_{j}\mid a_{j-1})$
11:    end for
12:    $\hat{p}\leftarrow\max(p,\varepsilon)$Ensure $\hat{p}\geq\varepsilon$}
13:    $\mathrm{ER}_{\text{sum}}\leftarrow\mathrm{ER}_{\text{sum}}-m_{i}\,\log_{2}(\hat{p})$$m_{i}$: multiplicity of $t_{i}$}
14:end for
15:$\overline{\mathrm{ER}}\leftarrow\mathrm{ER}_{\text{sum}}/T$Divide by total traces}
16:return  $\overline{\mathrm{ER}}$
Additionally, as mentioned above, the original ER formulation also handles non-fitting traces using a background model. However, in our method, ER scores are computed only for trace–DFG combinations in which the trace was used to discover the DFG. Thus, the model is guaranteed to fit the trace and its probability can be directly derived from the DFG without a background model. Finally, to avoid numerical underflow when computing probabilities for complex models (e.g., traces with many loops), the implementation constrains each trace probability to be at least $10^{-10}$ before taking the logarithm.
3.2 Entropic Clustering
Our method, called Entropic Clustering (EC), aims to minimize the entropic relevance (ER) scores of the cluster-DFG pairs (i.e. the ER score of the cluster of the DFG obtained from that cluster) within each cluster, as outlined in Algorithm 2. To improve efficiency, we make use of a variant log, consisting of the unique control-flow variants and their occurrence counts, rather than processing each individual case separately. The algorithm begins by specifying the required number of clusters, $k$, and selecting initial seed variants (explained below) to form the first clusters. A DFG is then discovered for each cluster. Next, we iterate through the remaining variants in order of prevalence, compute their ER scores with each cluster’s DFG, and assign them to the cluster where they achieve the lowest ER score. After assignment, that cluster’s DFG is updated by adjusting its node and edge counts accordingly. This assignment process is illustrated in Figure 1. The trace is hypothetically added to each cluster’s DFG, its ER score is computed, and it is assigned to the cluster where the ER increase is minimal, optimizing the DFG graph to have as few low frequency DFs as possible.
Since ER computation scales linearly with input size [5], the clustering process (excluding initialization) is also linear in complexity with the exact runtime depending on trace lengths, the preselected number of clusters, and the number of unique variants. An alternative approach, which adds the variant to the cluster for which the total cluster ER score is minimal (after adding), rather than the specific variant’s ER, was also explored. However, initial results showed inferior performance, leading us to omit it from further experiments.
Algorithm 2 Entropic Clustering (EC)
0: Event log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$, where $t_{i}$ is a trace variant and $m_{i}$ its multiplicity; number of clusters $k$
0: Partitioning of $L$ into $k$ clusters $\mathcal{C}=\{C_{1},\ldots,C_{k}\}$
1: Let $\mathcal{V}\leftarrow\{t_{1},\ldots,t_{K}\}$ be the set of unique variants in $L$
2: Select $k$ initial seeds $\{s_{1},\ldots,s_{k}\}\subseteq\mathcal{V}$ using a chosen initialization method
3:for $j=1$ to $k$ do
4:    Initialize cluster $C_{j}\leftarrow\{s_{j}\}$
5:    Build DFG $G_{j}$ from $s_{j}$
6:end for
7:for each variant $v\in\mathcal{V}\setminus\{s_{1},\ldots,s_{k}\}$, in decreasing order of multiplicity do
8:    for each cluster $C_{j}\in\mathcal{C}$ do
9:       Temporarily add $v$ to $C_{j}$ and update DFG $G_{j}$
10:       Compute average entropic relevance $\mathrm{ER}(v,G_{j})$
11:    end for
12:    Let $j^{*}=\arg\min_{j}\mathrm{ER}(v,G_{j})$
13:    Assign $v$ to cluster $C_{j^{*}}$ and update $G_{j^{*}}$ accordingly
14:end for
15:return  Final cluster assignment $\mathcal{C}=\{C_{1},\ldots,C_{k}\}$
Figure 1: Illustrative example how the method selects the cluster.
Next to this, we also implemented an alternative variant of EC that iteratively splits the cluster with the highest ER score until the required number of clusters is reached (Algorithm 3). The motivation behind this approach is that by continuously refining the worst-performing cluster, we may achieve overall better results, similar to decision trees employing ID3 or CART.
Algorithm 3 Entropic Clustering Split (EC-Split)
0: Event log $L=\{(t_{i},m_{i})\}_{i=1}^{K}$, desired number of clusters $k$
0: Partitioning of $L$ into $k$ clusters $\mathcal{C}=\{C_{1},\ldots,C_{k}\}$
1: Initialize clustering $\mathcal{C}\leftarrow\text{EC}(L,2)$Run Entropic Clustering with 2 clusters}
2:while $|\mathcal{C}|<k$ do
3:    for each cluster $C_{j}\in\mathcal{C}$ do
4:       Compute average entropic relevance $\overline{\mathrm{ER}}(C_{j})$
5:    end for
6:    Let $j^{*}=\arg\max_{j}\overline{\mathrm{ER}}(C_{j})$Identify worst-fitting cluster}
7:    Let $L_{j^{*}}$ be the variant log of cluster $C_{j^{*}}$
8:    Replace $C_{j^{*}}$ with two sub-clusters $\{C_{j^{*}1},C_{j^{*}2}\}\leftarrow\text{EC}(L_{j^{*}},2)$
9:    Update $\mathcal{C}\leftarrow(\mathcal{C}\setminus\{C_{j^{*}}\})\cup\{C_{j^{*}1},C_{j^{*}2}\}$
10:end while
11:return  Final cluster assignment $\mathcal{C}$
3.3 Initialization
The EC algorithm requires initial seed selection, which determines the first traces assigned to each cluster. The simplest approach is random selection, but since these initial seeds remain in their respective clusters, their selection can significantly impact clustering quality. To address this, we propose a more sophisticated ++ initialization (Algorithm 4), inspired by how k-means++ clustering handles initialization [22]. This method enhances seed selection by considering pairwise ER distances between traces. This initialization starts again by picking the first seed at random. Then, the pairwise ER distance between this seed and all other variants (traces) is computed. The pairwise ER is calculated by constructing a DFG from just these two traces, and calculating the average ER scores over both traces. The more similar those two traces are, the lower this score (analogous to edit distance). The next seed is selected randomly, weighted by the square of its pairwise ER distance from the current seed. We then continue similarly for the next seeds (if needed), but now calculate the pairwise ER with each seed for each variant (trace). For each variant, we use the minimum distance, i.e., the distance to the closest seed, as weights when selecting the next seed. We continue until we have a number of seeds equal to our required number of clusters.
Algorithm 4 ++ Initialization for Entropic Clustering
0: Set of trace variants $\mathcal{V}=\{t_{1},\ldots,t_{K}\}$, desired number of seeds $k$
0: Set of selected seeds $\mathcal{S}=\{s_{1},\ldots,s_{k}\}\subseteq\mathcal{V}$
1: Randomly select $s_{1}\in\mathcal{V}$
2: Initialize seed set $\mathcal{S}\leftarrow\{s_{1}\}$
3:while $|\mathcal{S}|<k$ do
4:    for each variant $v\in\mathcal{V}\setminus\mathcal{S}$ do
5:       for each seed $s\in\mathcal{S}$ do
6:          Construct DFG from $\{v,s\}$
7:          Compute average entropic relevance $\mathrm{ER}(v,s)$
8:       end for
9:       Let $d(v)=\min_{s\in\mathcal{S}}\mathrm{ER}(v,s)$
10:    end for
11:    Sample new seed $s^{*}\in\mathcal{V}\setminus\mathcal{S}$ with probability proportional to $d(v)^{2}$
12:    Update $\mathcal{S}\leftarrow\mathcal{S}\cup\{s^{*}\}$
13:end while
14:return  Final seed set $\mathcal{S}$
In the absence of any kind of looping behavior, the ER of a trace with its single-trace DFG, i.e. a DFG discovered by only itself, will be equal to 0. In this sense we mean any type of repetition of activity labels looping behavior. Without any form of repetition, the single-trace DFG will of course always simply be linear, with each edge probability equal to 1. However, when we have loops this is not necessarily the case, since at least one node (the repeated activity type) will have multiple outgoing edges, leading to edge probabilities not equal to 1. If a loop is taken multiple times, the (pairwise) ER score of that trace will be high, with any DFG. This is illustrated with an example in Figure 2. This effect can also have a significant impact on our initialization, cases containing loops will more likely be selected as initial seeds. To mitigate this, we also implemented a normalized version ++norm where we divide the components of the pairwise ER scores by the ER of that variant with its own single-trace DFG. This ensures that initialization is less biased towards selecting only traces with multiple loops or other types of repeating behavior, as seeds.
Figure 2: Illustrative example showing how loops can make the ER of a single-trace DFG non-zero.
4 Experimental setup
Our experimental setup evaluates the EC algorithm against various clustering methods by addressing the following key questions:
  • How uniformly structured are the clusters from a stochastic perspective?
  • How simple and readable are the discovered process models for each cluster?
  • What impact does the clustering method have on (non-stochastic) fitness and precision.
4.1 Event logs
In order to test our methods on a variety of different process behavior, we used 8 real-life event logs: the helpdesk event log [23], the road traffic fine management (RTFM) event log [24], the BPIC 2013 challenge event logs (the incidents and closed problems) [25], the hospital billing event log [26], the sepsis event log [27], and BPIC 2012 [28] and 2015 event logs [29]. Table 1 provides an overview of the key characteristics of these event logs.
Table 1: Log Information
Log Name
Total Cases
# Variants
Voc. Size
Avg. Case Len.
Min. Case Len.
Max. Case Len.
k
Helpdesk 4,580 226 14 4.66 2 15 4
RTFM 150,370 231 11 3.73 2 20 4
BPIC13 In. [^1] By using the default importer of pm4py activity types were not separated into lifecycle transitions. 7,554 1,511 4 8.68 1 123 5
BPIC13 Cl. [^1] By using the default importer of pm4py activity types were not separated into lifecycle transitions. 1,487 183 4 4.48 1 35 5
Hospital Billing 100,000 1,020 18 4.51 1 217 5
Sepsis 1,050 846 16 14.49 3 185 6
BPIC12 13,087 4,366 24 20.04 3 175 4
BPIC15 5,647 2,501 29 16.37 3 76 4
4.2 Evaluation metrics
Evaluating clustering methods is inherently challenging, as the quality of clusters depends on downstream tasks. To address this, and to stay close to the end-goal of most trace clustering pipelines in process mining, we employ multiple evaluation metrics related to the quality and readability of the discovered process models in each cluster. First of all, we use the Inductive Miner infrequent variant [8], with a (default) noise threshold of 0.2, to discover a Petri net for each cluster. Using the clustered event log and this Petri net we then calculate different (non-stochastic) conformance checking metrics: namely token-based fitness and precision [10], and alignment-based fitness [11] and precision [12]. Next to this, we also calculate the inverse arc-degree to measure the simplicity of the discovered Petri nets [13]. For these we used the implementation in the Python package pm4py [^1] Version 2.7.8.4 [30]. Next, we also discover a DFG (without filters) for each cluster and calculate the entropic relevance as explained before [5] using our simplified version as described above, as we only have fully fitting process models; providing us with a stochastic conformance checking measure. To reiterate: note that this is an average ER score over all traces in the event log, rather than a sum over all traces. To quantify the structural uncertainty of the graph, we also calculate the total graph entropy (GE) directly with only the DFG. Specifically, we compute the Shannon entropy of edge transitions using the probabilities of going from one node (activity) to another.
$GE=\sum_{u\in V}H_{u}=-\sum_{u\in V}\sum_{v\in V,p(u\to v)>0}p(u\to v)\log_{2}p(u\to v)$
Higher entropy indicates a more unpredictable structure, while lower entropy suggests a more deterministic or structured connectivity pattern [31]. Next to GE, we also introduce graph density (GD) as a measure on our DFGs, which measures how close the graph is to being fully connected. Graph density is calculated by dividing the number of edges actually present (non zero) in the DFG, by the maximum number of possible edges:
$GD=\frac{|\text{Edges}|}{|\text{Activity types}|\cdot(|\text{Activity types}|-1)}$
This provides us with another (non-stochastic) simplicity measure, but this time measuring the DFG’s connectivity.
A high-level overview of the experimental setup can be found in Figure 3.
Figure 3: High level overview of the experimental setup.
4.3 Baseline methods
To assess the performance of our proposed methods, we compare them with several baseline and existing clustering approaches. First, we include two simple baselines. The first baseline does not apply any clustering, meaning that the entire event log is used when calculating the evaluation metrics. The second baseline applies a random clustering approach, where traces are randomly assigned to clusters. Beyond these baselines, we also compare our methods with a frequency-based clustering approach. In this method, we construct a one-hot-encoded vector for each trace, where each dimension of the vector corresponds to an activity type and the values represent the frequency of occurrences of these activities. The vectors are then normalized by dividing each value by the total number of activities in the respective trace. Once the vectors are prepared, clustering is performed using the k-means++ algorithm [22]. We also compare our methods to an embedding-based clustering method using trace2vec [17]. Trace2vec is a self-supervised doc2vec-style approach [32], where trace embeddings are learned by training a shallow neural network to predict the context of activities within a sequence. In our implementation, we use a distributed bag-of-words (PV-DBOW) model trained for 200 epochs. The embedding dimension is set to the square root of the number of unique activity types, while the window size is set to two activities. We use the Gensim library [^2] Version 4.3.2 [33]. After obtaining the trace embeddings, we again apply the k-means++ algorithm to obtain clusters. We further compare our approach with the active trace clustering method ActiTraC, in both its frequency-based and distance-based variant. For both ActiTraC methods we use the implementation within the ProM framework [^3] Version 6.13 [34], and the default parameter settings, where the minimal cluster size is set to 0.25, the target ICS-fitness is set to 1.0, and the heuristic net discovery is used. To maintain consistency across all methods, the number of clusters is set to be the same as in the other approaches. Any traces that are not initially assigned to clusters are distributed among the existing clusters.
4.4 Elbow experiment
Since entropic clustering requires a predefined number of clusters, we conduct an elbow experiment to determine a reasonable cluster count. In this experiment, we plot the values of entropic relevance, graph density, and graph entropy across different numbers of clusters. We perform this analysis for the efficient random clustering, frequency-based clustering, EC, and ECsplit (both with ++ initialization). The results of these experiments can be found in the online repository. For some event logs, a clear elbow point with consensus among metrics and methods could be achieved, but for others this was more difficult. Nevertheless, since our downstream experiments focus on comparing different clustering methods under the same fixed number of clusters, a potentially arbitrary choice of cluster count does not significantly impact the experimental validity. The specific number of clusters selected for each event log is provided in Table 1.
4.5 Ranking
To evaluate the effectiveness of each clustering method, we determine its ranking for each event log along each metric. For each metric, we then calculate the average rank of each clustering method. The statistical significance of the ranking is determined following the well-established methodology described by Demšar et al. [35]. Specifically, for each metric, we perform Friedman’s test to assess whether the ranking of clustering methods is statistically significant across datasets. Friedman’s test is a non-parametric test that uses the average ranks of clustering methods across datasets. The null hypothesis assumes that all methods perform equivalently on average (i.e., their ranks should be equal). The Friedman test statistic follows a chi-square distribution with $\chi^{2}_{F}$ and $k-1$ degrees of freedom, where $k$ is the number of clustering methods. If $p<0.05$, we reject the null hypothesis, indicating that at least one method differs significantly. If Friedman’s test is significant, we proceed with post-hoc pairwise comparisons using the Nemenyi test, which determines whether the performance differences between clustering methods are statistically meaningful, and accounts for multiple comparisons. The results from the Nemenyi test are further visualized using a Critical Difference (CD) diagram, which provides an intuitive way to interpret the findings. In a CD diagram, clustering methods are placed along a ranking scale, and those that are not significantly different are connected with a horizontal line.
5 Experimental results
For each event log, the same number of clusters is extracted using each clustering method. The different metrics, as described above, were computed for each cluster, alongside the metric values for the full event log, which served as a baseline. The final score for each method is determined by taking the weighted average of the scores, where the weight is based on the number of cases in each cluster, ensuring that each case is weighted equally. The detailed results for each event log can be found in Appendix A. Due to the experiments running out of time the alignments-based conformance checking methods were not used for logs BPIC12 and BPIC15. The full results, for each separate cluster, can be found online. Table 2 presents the average rankings of each clustering method across the eight event logs, evaluated separately for each metric. The table uses a color-coded ranking system, where green represents the highest rank (1) and white represents the lowest (12). The best-performing method for each metric is highlighted in bold with an underline. The Friedman $\chi^{2}$ statistic and corresponding p-value indicate whether at least one method significantly outperforms the others for a given metric. Statistically significant p-values (i.e., below 0.05) are marked in bold, which in this case applies only to simplicity, ER, graph density, and entropy. Notably, the proposed EC methods outperform the rest on these metrics.
Based on the ER scores, we can conclude that EC clustering leads to the most clusters, whose DFGs contain the least infrequent DFs , which is expected given that it optimizes for this property. Metrics reflecting the simplicity of DFGs within each cluster, as well as the arc degree of Petri nets discovered using the inductive miner, also rank EC methods highest. The splitting variant of EC generally works worse than the regular EC method, which directly initializes the correct number of clusters, though there are some exceptions. Although statistical significance could not be established across all datasets, it is evident that EC methods tend to perform worse on precision-based metrics—particularly when compared to activity frequency-based clustering.
Table 2: Average ranking of each clustering method 1-12 (including no clustering), along each metric, over 8 event logs.
method
replayfit
replayprec
alignfit [^1] Methods alignfit and alignprec did not run on BPIC12 and BPIC15 sets (out-of-time).
alignprec [^1] Methods alignfit and alignprec did not run on BPIC12 and BPIC15 sets (out-of-time).
simplarc
ER
graphdens
graphentr
full log 7.25 6.75 6.17 5.50 11.50 12.00 11.00 12.00
random 8.17 8.33 8.88 8.62 9.00 10.33 8.67 9.67
frequencykm++ 4.62 3.88 5.33 2.83 6.38 8.25 6.12 4.62
trace2veckm++ 7.38 4.62 7.67 4.33 6.00 9.12 8.25 5.88
actitracdist 7.88 5.75 5.50 6.00 6.62 7.88 8.62 7.38
actitracfreq 7.62 8.25 6.83 7.33 8.50 8.50 8.38 8.75
EC++ 4.62 7.75 6.67 7.67 5.50 2.00 4.38 4.38
EC++norm 6.25 7.38 6.83 8.17 4.25 2.12 3.38 3.88
ECrand 6.12 6.75 5.33 7.50 3.62 2.88 2.88 4.50
ECsplit++ 6.88 6.25 8.17 6.17 5.25 6.00 6.25 7.00
ECsplit++norm 6.25 6.00 6.17 7.00 5.75 4.38 5.62 5.62
ECsplitrand 4.75 6.75 4.83 7.33 6.25 4.38 4.62 4.00
Friedman $\chi^{2}$ 11.38 11.54 6.97 13.41 30.35 76.25 41.38 46.10
p-value 0.412 0.399 0.801 0.267 0.001 0.000 0.000 0.000
Figures 4(a) to 4(d) display the critical difference (CD) diagrams for simplicity (of discovered Petri nets), as well as the ER, graph density, and graph entropy of the corresponding DFGs. Notably, many methods are connected in the diagrams, meaning their differences are not statistically significant. In fact, most benchmark clustering methods do not perform significantly better than random clustering or no clustering at all. However, the best-performing EC variants demonstrate a statistically significant advantage in ER and graph entropy compared to these naive approaches. Since the different initialization techniques for EC clustering are all connected in the diagrams, we cannot conclude that any one approach is significantly superior to the others. This could indicate that the methods are less initialization-sensitive than initially thought.
(a) Critical difference diagram for the metric arc degree simplicity.
(b) Critical difference diagram for the metric ER.
(c) Critical difference diagram for the metric graph density.
(d) Critical difference diagram for the metric graph entropy.
Figure 4: Critical difference diagrams for different metrics.
6 Discussion & limitations
The experimental results indicate that our entropic clustering approaches lead to more structured clusters when taking a stochastic viewpoint over the log (i.e. whose underlying DFGs contain fewer infrequent DFs), as evidenced by lower ER scores. This, in turn, generally results in simpler downstream process models, both in terms of directly-follows graphs (DFGs) and Petri nets discovered using the Inductive Miner. It is, however, impossible to draw general conclusions about the quality of clustering methods concerning the fitness of the discovered Petri nets. In general, clustering based on activity frequency and trace2vec tends to produce more precise models, than our EC-based methods. These lower precision scores on the discovered Petri nets could perhaps be attributed to their optimization for DFGs, only focusing on DFs and therefore potentially generalizing too much behavior, potentially overlooking aspects such as parallelism. Since conformance checking in our experiments was performed on models mined using the Inductive Miner, the power and limitations of this discovery algorithm also influence the results.
In our current experimental setup, we did not include other stochastic conformance checking metrics. Other conformance checking metrics, such as using the Earth mover’s distance [14], would be most meaningful if we were not using fully fitting (non-filtered) DFGs as the discovered model. We would therefore have to decide on one stochastic discovery technique, which could have a large influence on the end results. Or, to properly account for this, try different techniques (and filter) leading to exponentially more experiments. We opted for ER as our primary measure because it captures the (stochastic) the lack of structure in clusters in a general way, independent of the possible downstream discovery.
While statistical ranking tests provide a more robust evaluation of different methods compared to simply reporting scores per event log, they do not indicate which method works best for specific types of processes. Another limitation of ranking-based evaluation is that if all methods perform poorly for a particular log, the ranking becomes somewhat arbitrary.
Unlike methods such as k-means, our approach does not rely on shifting cluster centroids to reach an equilibrium, as defining efficient cluster centers or medoids for process behavior is not straightforward. Consequently, the EC method could still remain somewhat sensitive to initialization in its current form. Our ++ and ++norm initializations still select the first seed randomly, which can lead to suboptimal starting conditions, and for the recursive ECsplit variant at each split. A more computationally expensive approach, incorporating all pairwise distances, could be explored to mitigate this issue. Additionally, the ++ and ++norm initializations, despite the latter’s normalization to account for loops, tend to favor outlier traces as initial seeds. While this may be beneficial in some cases, it could be suboptimal for certain applications. The lack of statistically significant differences between initialization methods may point to a lower-than-expected sensitivity to initialization. One might consider using the sum of all trace ER scores per cluster, rather than the average, to better reflect the the actual information compression cost. However, this choice does not affect any of the conclusions of this work. For our clustering algorithm, the impact would be negligible, as it relies solely on the ER scores of candidate traces. While the evaluation scores might differ slightly due to this change, we employ weighted averages, ensuring that the relative ranking of the algorithms remains unchanged.
7 Conclusion & future work
In this work, we introduced a trace clustering paradigm focused on reducing the complexity of underlying clusters from a stochastic process model perspective. Our entropic clustering techniques actively utilize the directly-follows graph (DFG) within each cluster, assigning traces by minimizing entropic relevance. Through extensive experiments on real-life event logs, we demonstrated that our methods outperform alternatives in minimizing entropic relevance while also producing less complex downstream process models, both in terms of DFGs and discovered Petri nets. Additionally, our experiments highlight how the perception of clustering quality changes depending on whether stochasticity is taken into account. Our approach is highly efficient, optimizing for commonly used DFGs, making it a viable candidate for integration into commercial process mining tools.
For future work, a larger-scale experiment could be conducted, potentially using controlled synthetic data. This would allow us to systematically analyze which types of process behaviors and challenges affect different clustering techniques. A simple but useful extension could involve introducing a robust version of the algorithm that detects and handles outliers, as they likely skew clustering results in the current approach. Further extensions could expand the model-driven stochastic clustering paradigm by incorporating other process model types, discovery algorithms, and stochastic conformance checking techniques. We opted for the current combination due to its efficiency and widespread adoption in commercial tools. Our algorithm (see Alg. 2) updates each cluster’s DFG incrementally when assigning traces. An alternative approach could forgo this update and use the DFGs as they are, relying on an ER metric implementation [5] including an appropriate background model for non-fitting traces. Another avenue for improvement lies in the initialization phase. Seed selection could be guided by expert knowledge, if available. Given the scalability of our method, running the clustering multiple times and selecting the best result could help mitigate volatility due to initialization. Additionally, since determining the correct number of clusters upfront is often impractical, a variant that dynamically adjusts the number of clusters, based on a maximum ER threshold, for example, could be explored. Finally, our approach could be extended to incorporate additional process features, such as the resource performing an activity or sensor measurements. One way to achieve this would be to adjust probability calculations based on these features. For categorical attributes, probabilities could be derived from frequency distributions in the current traces, while for continuous attributes, they could be estimated using a fitted (normal) distribution.
Acknowledgments
This work was supported in part by the Research Foundation Flanders (FWO) under Project 1294325N as well as grant number G039923N, and Internal Funds KU Leuven under grant number C14/23/031.
Appendix A Results for the Individual Event Logs
In this appendix we present the results for each event log separately. Columns correspond to the different metrics and rows with the different clustering techniques. We report weighted averages, i.e., to calculate the metric value for a clustering the measurements for each cluster are weighted by the amount of cases in that cluster.
Table 3: Weighted averages Helpdesk
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.989 0.781 0.978 0.781 0.682 3.618 0.279 12.729
random 0.992 0.695 0.984 0.695 0.690 2.754 0.268 7.912
frequencykm++ 0.995 0.831 0.990 0.831 0.686 2.670 0.259 4.779
trace2veckm++ 0.983 0.845 0.977 0.845 0.696 2.744 0.243 6.367
actitracdist 0.993 0.659 0.986 0.659 0.682 2.434 0.253 7.743
actitracfreq 0.919 0.885 0.918 0.886 0.686 3.026 0.266 8.807
EC++ 0.974 0.607 0.973 0.607 0.717 2.108 0.194 5.846
EC++norm 0.974 0.596 0.973 0.596 0.716 2.105 0.198 5.839
ECrand 0.969 0.664 0.968 0.664 0.699 2.118 0.216 6.094
ECsplit++ 0.967 0.750 0.954 0.750 0.702 2.379 0.276 7.241
ECsplit++norm 0.971 0.757 0.961 0.757 0.698 2.387 0.292 7.152
ECsplitrand 0.987 0.757 0.977 0.757 0.692 2.482 0.292 7.037
Table 4: Weighted averages RTFM
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.929 0.813 0.878 0.813 0.630 2.935 0.500 11.428
random 0.978 0.660 0.971 0.660 0.676 1.453 0.419 10.687
frequencykm++ 0.997 0.880 0.993 0.880 0.693 1.708 0.283 7.407
trace2veckm++ 0.991 0.925 0.987 0.925 0.746 1.229 0.341 9.038
actitracdist 0.994 0.825 0.992 0.825 0.768 0.961 0.369 5.157
actitracfreq 0.968 0.752 0.955 0.752 0.681 1.250 0.403 8.415
EC++ 0.995 0.818 0.988 0.818 0.776 0.730 0.333 8.567
EC++norm 0.946 0.682 0.950 0.682 0.726 0.730 0.300 7.720
ECrand 0.995 0.792 0.988 0.792 0.774 0.731 0.326 8.425
ECsplit++ 0.981 0.735 0.980 0.735 0.742 0.960 0.388 7.756
ECsplit++norm 0.981 0.687 0.980 0.687 0.676 0.960 0.366 7.780
ECsplitrand 0.981 0.665 0.980 0.665 0.707 0.958 0.355 7.733
Table 5: Weighted averages Hospital billing
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.990 0.591 0.975 0.591 0.629 3.450 0.416 18.223
random 0.993 0.463 0.979 0.463 0.639 2.307 0.318 15.437
frequencykm++ 0.995 0.637 0.987 0.637 0.648 2.156 0.258 11.383
trace2veckm++ 0.984 0.472 0.964 0.472 0.635 2.410 0.275 10.290
actitracdist 0.988 0.787 0.957 0.787 0.685 1.942 0.337 11.546
actitracfreq 0.991 0.627 0.988 0.627 0.631 1.968 0.304 12.834
EC++ 0.998 0.615 0.995 0.618 0.658 1.237 0.249 13.752
EC++norm 0.995 0.691 0.985 0.697 0.670 1.232 0.240 13.214
ECrand 0.995 0.674 0.984 0.677 0.671 1.239 0.239 13.512
ECsplit++ 0.988 0.669 0.967 0.669 0.666 1.356 0.309 14.327
ECsplit++norm 0.983 0.642 0.963 0.642 0.663 1.346 0.237 11.695
ECsplitrand 0.984 0.701 0.963 0.702 0.657 1.355 0.230 11.167
Table 6: Weighted averages Sepsis
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.967 0.619 0.962 0.619 0.593 24.444 0.441 23.580
random 0.954 0.663 0.949 0.679 0.601 23.983 0.364 21.753
frequencykm++ 0.978 0.672 0.962 0.674 0.614 22.594 0.342 18.925
trace2veckm++ 0.972 0.602 0.956 0.602 0.629 23.175 0.499 18.212
actitracdist 0.968 0.567 0.967 0.575 0.606 22.695 0.367 20.646
actitracfreq 0.976 0.610 0.969 0.610 0.623 22.621 0.379 20.710
EC++ 0.961 0.532 0.948 0.536 0.601 21.163 0.326 18.841
EC++norm 0.974 0.539 0.958 0.543 0.603 21.324 0.327 18.977
ECrand 0.979 0.554 0.960 0.558 0.610 21.220 0.317 18.697
ECsplit++ 0.964 0.506 0.953 0.512 0.608 22.243 0.338 19.563
ECsplit++norm 0.971 0.486 0.959 0.490 0.596 22.039 0.325 19.497
ECsplitrand 0.986 0.455 0.971 0.457 0.612 22.040 0.332 19.332
Table 7: Weighted averages BPIC13 incidents
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.966 0.841 0.957 0.841 0.733 10.728 0.533 3.668
random 0.959 0.866 0.953 0.866 0.734 10.498 0.586 3.558
frequencykm++ 0.987 0.947 0.930 0.960 0.770 9.704 0.522 3.079
trace2veckm++ 0.945 0.877 0.954 0.877 0.697 10.452 0.653 3.657
actitracdist 0.988 0.844 0.983 0.844 0.722 9.865 0.528 3.287
actitracfreq 0.988 0.775 0.983 0.775 0.722 9.867 0.516 3.305
EC++ 0.995 0.839 0.950 0.854 0.766 9.191 0.540 2.879
EC++norm 0.995 0.866 0.986 0.866 0.784 9.191 0.455 2.544
ECrand 0.991 0.833 0.979 0.833 0.771 9.445 0.513 2.953
ECsplit++ 0.997 0.909 0.964 0.915 0.769 9.767 0.537 3.319
ECsplit++norm 0.997 0.926 0.991 0.926 0.772 9.506 0.561 3.143
ECsplitrand 0.997 0.849 0.994 0.849 0.748 9.380 0.504 3.037
Table 8: Weighted averages BPIC13 closed problems
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.997 0.971 0.990 0.971 0.750 5.047 0.500 3.916
random 0.989 0.802 0.957 0.802 0.766 4.675 0.380 2.455
frequencykm++ 0.986 0.958 0.969 0.959 0.783 4.075 0.452 2.346
trace2veckm++ 0.993 0.941 0.977 0.941 0.776 4.717 0.448 2.711
actitracdist 0.987 0.901 0.970 0.901 0.775 4.342 0.453 2.819
actitracfreq 0.986 0.796 0.966 0.796 0.770 4.567 0.458 3.154
EC++ 0.997 0.907 0.989 0.907 0.772 3.826 0.371 1.622
EC++norm 0.996 0.813 0.989 0.813 0.774 3.792 0.337 1.625
ECrand 0.996 0.850 0.990 0.850 0.773 3.798 0.344 1.953
ECsplit++ 0.996 0.938 0.989 0.938 0.758 4.022 0.394 2.751
ECsplit++norm 0.998 0.813 0.994 0.813 0.768 3.921 0.407 2.782
ECsplitrand 0.996 0.922 0.990 0.922 0.764 3.975 0.399 2.588
Table 9: Weighted averages BPIC12
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.999 0.150 / / 0.600 19.599 0.211 34.259
random 0.971 0.342 / / 0.625 18.803 0.204 32.639
frequencykm++ 0.973 0.532 / / 0.665 18.626 0.189 30.663
trace2veckm++ 0.971 0.535 / / 0.686 18.579 0.172 22.166
actitracdist 0.944 0.397 / / 0.652 17.326 0.191 27.323
actitracfreq 0.997 0.294 / / 0.643 16.963 0.176 26.933
EC++ 0.993 0.386 / / 0.640 16.798 0.171 24.821
EC++norm 0.993 0.369 / / 0.658 16.861 0.163 24.149
ECrand 0.964 0.680 / / 0.690 16.841 0.140 19.965
ECsplit++ 0.966 0.690 / / 0.714 17.030 0.151 20.643
ECsplit++norm 0.965 0.622 / / 0.699 16.980 0.151 20.009
ECsplitrand 0.969 0.624 / / 0.707 16.912 0.151 18.398
Table 10: Weighted averages BPIC15
method replayfit replayprec alignfit alignprec simplarc ER graphdens graphentr
full log 0.884 0.366 / / 0.572 26.273 0.308 29.932
random 0.901 0.415 / / 0.625 26.123 0.235 26.660
frequencykm++ 0.938 0.382 / / 0.600 24.584 0.235 24.045
trace2veckm++ 0.938 0.397 / / 0.612 24.351 0.246 22.264
actitracdist 0.905 0.401 / / 0.612 25.223 0.245 25.279
actitracfreq 0.930 0.364 / / 0.606 25.146 0.244 24.257
EC++ 0.937 0.397 / / 0.620 24.250 0.223 20.707
EC++norm 0.916 0.469 / / 0.625 24.387 0.240 22.087
ECrand 0.930 0.395 / / 0.623 24.330 0.239 22.198
ECsplit++ 0.943 0.354 / / 0.614 24.393 0.218 21.741
ECsplit++norm 0.940 0.432 / / 0.642 24.191 0.223 20.883
ECsplitrand 0.942 0.384 / / 0.616 24.249 0.217 20.916